Integrate the following functions w.r.t. x:

1. 1/(2x+1)3/2

2. Sin(2x+3)

3. Cosec(4x)

4. 1/(1-9x2)1/2




1
Expert's answer
2021-06-07T17:22:56-0400

"\displaystyle
1)\\ \int \frac{1}{(2x+1)^{\frac{3}{2}}}\,\,\mathrm{d}x = -2(2x+1)^{\frac{-1}{2}} \times \frac{1}{2} + C = -(2x+1)^{\frac{-1}{2}} + C\\


2)\\ \int \sin(2x + 3)\,\,\mathrm{d}x = -\frac{\cos(2x + 3)}{2} + C\\

3)\\
\begin{aligned}
\int \cosec(4x)\,\mathrm{d}x &=
\int \cosec(4x)\frac{\cosec(4x) + \cot(4x)}{\cosec(4x) + \cot(4x)}\,\mathrm{d}x 
\\&= \int\frac{\cosec^2(4x) + \cosec(4x)\cot(4x)}{\cosec(4x) + \cot(4x)}\,\mathrm{d}x 
\\&= -\int\frac{\mathrm{d}\left(\cosec(4x) + \cot(4x)\right)}{4(\cosec(4x) + \cot(4x))}
\\&= -\frac{1}{4}\ln\left|\cosec(4x) + \cot(4x)\right| + C
\end{aligned}
\\

4)\\ \begin{aligned}
\int \frac{\mathrm{d}x}{\sqrt{1 - 9x^2}} &= \frac{1}{\sqrt{9}}\int \frac{\mathrm{d}x}{\sqrt{\frac{1}{9} - x^2}} 
\\&= \frac{1}{3}\int \frac{\mathrm{d}x}{\sqrt{\frac{1}{9} - x^2}} 
\\&= \frac{1}{3}\arcsin\left(\frac{x}{1/3}\right) + C 
\\&= \frac{1}{3}\arcsin\left(3x\right) + C
\end{aligned}"